Safawi kicked a ball with an initial speed of 20 ms-1 at 30o angle. If the ball experienced a constant vertical acceleration of -9.81 ms-2, calculate the ball’s maximum height and distance.


1
Expert's answer
2022-02-08T10:12:26-0500

"h=\frac{v^2\sin^2\alpha}{2g}=5.1~m,"

"l=\frac{v^2\sin2\alpha}g=35.3~m."